A Chinese restaurant makes a special sauce for its dishes which consists of oyster sauce and chili oil. Oyster sauce costs $5.96 per quart and chili oil costs $8.04 per quart. How much of each should the restaurant use to make 60 quarts of the special sauce if a price of $7.52 is desired?
Answer by checkley71(8403)

5.96X+8.04(60-X)=7.52*60 X BEING THE AMOUNT OF OYSTER SAUCE
5.96X+482.4-8.04X=451.2 OR -2.08X=-31.2 OR X=31.2/2.08 OR X=15 QUARTS OF OYSTER SAUCE THEREFORE THERE WERE 60-15=45 QUARTS OF CHILI SAUCE.